Colin, Dave and Emma share some money.
Colin gets 3/10 of the money.
Emma and Dave share the rest of the money in the ratio 3 : 2.
What is Dave's share of the money?

Answer by Boreal(15235)

Emma and Dave share 70%
In a 3:2 ratio, they apportion every $5 by giving Emma $3 and Dave $2. 70/5=14
14*3=$42 for Emily
14*2=$28 for Dave ANSWER